A primary standard of KHP is made up by dissolving 4.850g of KHP in a total volume of 250mL. An aliquot of 25.00mL of the standardised KHP is then titrated with a sodium hydroxide solution of unknown concentration. An average titre of 16.85mL of the sodium hydroxide was used to achieve the equivalence point. Calculate the concentration of the sodium hydroxide.
NB. KHP is a monoprotic acid.
1
Expert's answer
2020-08-02T15:52:03-0400

Solution:

KHP formula is C8H5O4K.

1.n = m/M = 4.850/204 = 0.02377 mol.

2.c = n/V = 0.02377/0.25 = 0.09509 mol/dm3.

3.c2 = c1V1/V2 = 0.09509*0.025/0.01685 = 0.141 mol/dm3.

Answer: 0.141 mol/dm3.